My sister was a quarter of my age when i was 14 years old. She is now 21 years old; How hold i am?
Answer by Theo(13342)

when you were 14 years old, your sister was 1/4 your age.
that makes her 14/4 = 3.5 years old when you were 14 years old.

she is now 21 years old.
that makes her 21 - 3.5 = 17.5 years older than when she was 3.5 years old.
that makes you 14 + 17.5 = 31.5 years old now.

your solution is that you are now 31.5 years old.